# Extracting transport coefficients from local ground-state currents

### v1

@@ Abstract
Transport properties are central to characterizing quantum matter, yet their extraction typically requires external forcing and time-resolved measurements. In this work, we propose a scheme to access transport coefficients directly from measurements of local, static ground-state currents – quantities readily accessible in quantum-engineered platforms. By exploiting the exponential decay of correlations in gapped systems and the finite velocity of correlation spreading, we demonstrate that the local Hall response can be reconstructed from a small set of quasi-local current observables. We derive explicit relations connecting these static observables to a practical local Chern marker, and introduce a scalable digital protocol for measuring the required generalized currents in cold-atom quantum simulators. Numerical simulations of a non-interacting Chern insulator validate our approach. Moreover, the scheme extends naturally to fractional Chern insulators and other strongly correlated systems, even at finite temperature, offering a broadly applicable route to probing transport in engineered quantum matter.
I Introduction
@@
Transport measurements are essential for identifying and characterizing exotic quantum matter, including superconducting states, strange metals, and topological insulators. While transport is traditionally probed by measuring currents upon connecting reservoirs to a sample, transport coefficients can also be extracted indirectly through other observables and protocols. Examples of such approaches include the density response to a magnetic perturbation following Středa’s formula [1, 2], circular dichroism via Kramers-Kronig relations [3, 4, 5] and related spectroscopic probes [6], as well as center-of-mass drifts upon acting with an external force [7, 8, 9]. These methods have proven to be powerful methods to probe transport coefficients in various settings, including synthetic quantum matter of cold atoms in continuous systems [10, 11] and optical lattices [12, 13, 14, 15, 16], as well as photonic systems [17, 18].
@@
Theoretically, the conductivity can be expressed as a current-current correlation function via the Kubo formula [19], suggesting that transport could be monitored by evaluating local currents in the ground state. As further discussed below, this would a priori require monitoring these currents over time, within a region located around a reference site delimited by the correlation length. While challenging in traditional solid-state platforms, this is particularly appealing to recent quantum-engineered setups, where local currents can be measured with great accuracy [20]. Local transport properties can thus be extracted, which, for topological insulating states, provides a practical scheme to access local probes of topology, e.g. local Chern markers [21, 22, 23]. At this stage, however, the measurement of temporal correlations is experimentally challenging; see Ref. [24] for a proposal based on many-body Ramsey interferometry.
@@
In this work, we introduce a scheme to extract the local Hall response (Chern marker) of a gapped many-body system, which only requires local ground-state current measurements. In contrast to more conventional approaches, this scheme entirely builds on static measurements, performed in the absence of any external perturbation. Importantly, this scheme exploits the exponential decay of correlations [25] in gapped systems, together with the finite velocity of the spread of correlations [26]. Under these conditions, the relevant information can be accessed from a power series of the time-evolved current-current correlations. As a consequence, the problem reduces to measuring equal-time current correlations within a finite region surrounding the reference site. Focusing on the transverse (Hall) conductivity, we demonstrate that the associated current–current correlator admits a substantial simplification: it can be expressed in terms of a limited set of further-range current observables. Crucially, we find that static ground-state current measurements alone suffice to accurately determine the Hall conductivity. This paradigm is summarized schematically in Fig. 1(a). Inspired by recent experiments in cold-atom quantum simulators [20], we provide a scalable, digital protocol to access the required further-range currents. We also present a numerical validation of our scheme, considering a non-interacting Chern insulator in the Hofstadter model.
Figure 1: (a) We propose to measure ground-state currents (red arrows) within a finite radius (gray shading) around a reference site (blue dot) to extract the local Hall response (Chern marker). (b) In a Chern insulator state, the current-current correlator $\mathcal{C}(\vec{r},t)$ defined in Eq. (4), relative to a reference site $\vec{r}$, exhibits damped oscillations, which are well captured by the ansatz in Eq. (6). (c) Upon fitting the damped oscillations, one can use the expression in Eq. (7) to obtain a local Chern marker. The result is in agreement with the Chern number ${\sf Ch}=1$ of the populated band. Here, the calculations were performed by filling the lowest band of the Hofstadter model with non-interacting fermions, at flux $\alpha=\nicefrac{{1}}{{q}}$ per plaquette.
@@
Since the correlation structures used in our protocol are common to all gapped, topological states, we expect our protocol to apply to a broad class of systems, such as strongly-interacting fractional Chern insulators or systems at finite temperature. Furthermore, similar protocols can be developed for other correlation functions, giving access to properties not easily accessible in traditional solid-state platforms.
II From the Hall conductivity to local ground-state currents
@@
Using Kubo’s formula, the local transverse conductivity, evaluated at position $\vec{r}$, reads
@@
II.1 The single-frequency ansatz
The unequal time correlation function contained in Eq. (4) cannot be accessed in current cold-atom quantum simulators. For gapped states, we can circumvent this limitation by approximating the correlator $\mathcal{C}(\vec{r},t)$ with an exponentially damped oscillation,
@@ II.1 The single-frequency ansatz
This approximation is applicable under the assumption that there is a single characteristic gap $\omega_{0}$, from the ground state to the relevant low-energy excitations. The rate $\Gamma$ then characterizes the broadening of this transition due to band dispersion and finite-size effects. This single-frequency ansatz is expected to hold in the regime of relatively flat Bloch bands, as we explicitly demonstrate below and in Appendix A.
@@
This expression offers an explicit recipe to obtain the local Hall response from the three parameters entering the single-frequency ansatz in Eq. (6): $\mathcal{C}(\vec{r},0)$, $\omega_{0}$ and $\Gamma$. Since the Hall response is directly related to the Chern number [29, 30], we point out that Eq. (7) provides a simple and practical expression for the local Chern marker ${\sf Ch}(\vec{r})$.
To confirm the validity of this approximation, we performed a numerical study of a paradigmatic Chern insulator state, which is obtained by filling the lowest band of the Hofstadter model with non-interacting fermions; see also the next Section III for details. We find that the correlator $\mathcal{C}(\vec{r},t)$ that was obtained numerically is in good agreement with the ansatz in Eq. (7); see Fig. 1(b). Here, we extracted the parameters $\Gamma$ and $\omega_{0}$ by fitting the numerical data. As shown in Appendix A, these parameters match the bandwidth $\Gamma\!\approx\!W$ and cyclotron gap $\omega_{0}\!\approx\!\Omega_{c}$ of the corresponding Hofstadter bands, respectively.
Furthermore, we confirm that the local Chern marker ${\sf Ch}(\vec{r})=2\pi\sigma_{\rm H}(\vec{r})$ [Eq. (7)], extracted from the fitted curve, agrees well with the expected Chern number ${\sf Ch}=1$ of the lowest Hofstadter band in the limit of sufficiently small $\alpha$; see Fig. 1(c). Deviations at large $\alpha$ are attributed to the stronger dispersion of the Hofstadter bands in this regime, which broadens the transition.
We explicitly confirm the local nature of the Chern marker ${\sf Ch}(\vec{r})$ in Appendix B, where we present the results for a junction displaying two (local and distinct) Chern insulating phases.
II.2 Static ground-state observables
Both in quantum-simulation experiments and in numerical studies of strongly correlated systems, it is challenging to extract the time-dependent current-current correlator $\braket{\hat{j}_{x}(\vec{r},t)\hat{j}_{y}(\vec{r}^{\prime},0)}$ in Eq. (4). However, the formal expression for this correlator can be rewritten in a more appealing form using the Baker-Campbell-Hausdorff identity,
@@ II.2 Static ground-state observables
Importantly, we remark that $c_{1}$ scales with the damping $\Gamma$, hence it is small whenever band-dispersion effects are limited. In fact, in our numerical investigations detailed below, we find that $c_{1}$ is indeed negligible, as well as higher odd corrections ($c_{3,5,\ldots}$). In contrast, we find that including even corrections ($c_{4,6,\ldots}$) improves the determination of the correlator $\mathcal{C}(\vec{r},t)$; see Appendix C. However, including those higher-order corrections leads to a significant overhead due to the numerous currents that need to be evaluated.
Furthermore, we remind that the coefficient $\omega_{0}$, which requires the measurement of $c_{2}$, corresponds to the characteristic gap in the flat-band regime. Thus, this coefficient could be evaluated independently through conventional spectroscopy. Altogether, this indicates that the measurement of the transport coefficient ${\sf Ch}(\vec{r})=2\pi\sigma_{\rm H}(\vec{r})$ can be limited to the evaluation of the single coefficient $c_{0}$ in the flat-band regime (i.e. the low-flux limit in the Hofstadter model).
Since the explicit form of the coefficients $c_{m}$ depends on the details of the Hamiltonian, we will now investigate a concrete example in the next Section, based on the Hofstadter model. However, we expect our approach to apply to a broad class of gapped, local Hamiltonians.
III Application to Chern insulators in the Hofstadter model
We consider non-interacting fermions in the Hofstadter model, defined on a square lattice, with flux $\alpha$ per plaquette. In the Landau gauge, the Hamiltonian reads
@@ III Application to Chern insulators in the Hofstadter model
where $\vec{r}=(x,y)$ denote the lattice sites and $J$ is the hopping amplitude. This well-known model exhibits topological bands [29], with Chern number ${\sf Ch}=1$ in the lowest band for $\alpha\!=\!1/q$, with $q$ an integer. The continuum limit of the model is reached upon setting $\alpha\rightarrow 0$, where the Hofstadter bands reduce to (flat) Landau levels.
III.1 The $c_{m}$ coefficients in terms of local currents
Figure 2: (a-c) Currents $\hat{j}_{\vec{\mu}}(\vec{R},\Theta)$ to be evaluated at (a) zeroth, (b) first, and (c) second order in $t$ for a Hofstadter model on a square lattice of $13\times 13$ sites with the reference site (blue) chosen in the center of the system. Darker arrows indicate that more than one generalized current with different phases $\Theta$ have to be measured. The necessary currents are found by first applying the iterative construction in Eq. (16), then extracting the imaginary part of the correlator using Eq. (18), and integrating over all lattice sites $\vec{R}$. (d) Local Chern marker ${\sf Ch}(\vec{r})=2\pi\sigma_{\rm H}(\vec{r})$ [Eq. (7)] as obtained from the coefficients $c_{m}$ in Eq. (10) using the experimental protocol. Here, the coefficients $c_{m}$ are connected to the parameters of Eq. (6) via Eq. (12). The values for the Chern marker obtained in this manner are in good agreement among each other as well as with the expected ${\sf Ch}=1$ for the lowest Hofstadter band. For additional details see Appendix C.
@@ III.1 The $c_{m}$ coefficients in terms of local currents
$\hat{j}_{\vec{\mu}}(\vec{r};\theta)=iJ\mathrm{e}^{i\theta}\hat{c}^{\dagger}_{\vec{r}+\vec{\mu}}\hat{c}^{\vphantom{\dagger}}_{\vec{r}}+\mathrm{H.c.}$ (15)
@@
We note that the number of summands in Eq. (19) is given by $8^{m}$ [Eq. (16)], with each summand contributing $4$ generalized currents [Eq. (18)]. Therefore, at $m$-th order $4\times 8^{m}$ currents need to be measured, in total resulting in $4\times(8^{0}+8^{1}+8^{2})=292$ current measurements to access the local Hall conductivity from the leading-order coefficients $c_{0,1,2}$. In Fig. 2(a-c) we show the currents that have to be evaluated for a Hofstadter model on a $13\times 13$ square lattice with the reference site chosen at the center of the system.
@@
We close this Section by reminding that the measurement of the transport coefficient ${\sf Ch}(\vec{r})=2\pi\sigma_{\rm H}(\vec{r})$ can be limited to the evaluation of the single coefficient $c_{0}$, in the flat-band regime; see Section II. This measurement would only require the determination of the four local diagonal currents displayed in Fig. 2(a), which is experimentally appealing.
III.2 Numerical study
@@
Next, we provide numerical evidence for the applicability of our protocol to the Hofstadter model. In our simulations, we consider an open square of size $L\times L=13\times 13$ with varying flux $\alpha=\nicefrac{{1}}{{q}}$, $q=3,\ldots,15$, per plaquette, so that the lowest band has Chern number ${\sf Ch}=1$. We construct a Chern insulator by occupying the lowest $N=\nicefrac{{(L_{x}\times L_{y})}}{{q}}$ states with fermions and use this state to test our experimental protocol. In particular, we implement the recursive generation of the currents visualized in Fig. 2(a-c) and evaluate their expectation values to obtain $c_{0,1,2}$. From these, we can extract the parameters of the damped oscillation [Fig. 1(b)] through Eq. (12) and insert them into the expression for the local Chern marker ${\sf Ch}(\vec{r})$ in Eq. (7).
@@
We find that the results obtained in this manner are in good agreement with the expected Chern number ${\sf Ch}=1$ of the lowest Hofstadter band, see Fig. 2(d). We emphasize again, that in this approach the problem reduced to the evaluation of $292$ generalized ground-state currents, instead of performing explicit time-evolution. While including the higher order contribution $c_{4}$ further improves the results, we find that neglecting the odd orders ($c_{1}$ and $c_{3}$) does not affect the results; see also Appendix C. We also numerically explore the applicability of our method to a fractional Chern insulator state in Appendix D.
III.3 Protocol for measuring further-range currents
Figure 3: Upper panel: Sketch of the pulse sequence needed to evaluate further-range currents. Lower left: Ground state density (dots) and $m=0$ currents (red arrows) of the Chern insulator in the lowest band of a Hofstadter model at $\alpha=\nicefrac{{1}}{{5}}$. The blue arrow indicates a possible path to measure a current between next-nearest neighbors. Lower right: Density difference after the pulse sequence and coefficients $\gamma_{i}$ with which the respective points along the path are weighted.
@@ III.3 Protocol for measuring further-range currents
The experimental protocol presented above requires the measurement of generalized, further-range currents connecting lattice sites $\vec{R}$ and $\vec{R}+\vec{\mu}$. Here, we present a scalable, digital protocol to achieve such measurements using only nearest-neighbor tunnelings $\hat{T}_{ij}$ and local energy offsets $\hat{M}_{i}$ on the lattice,
@@
and local density measurements on the sites involved.
@@
  • $\ell=3$: Here, we need a pulse of duration $t_{2}=\frac{\pi}{2\Delta}$ to obtain the required phase (up to a global sign). Here, the initial ground-state densities along the path contribute with $\gamma_{1}=\gamma_{3}=\frac{1}{2}$, $\gamma_{2}=-1$, see Fig. 3.
@@
IV Conclusion and Outlook
In this work, we showed that measurements of ground-state currents of gapped many-body systems are sufficient to extract the local Hall marker. In particular, using the quasi-local nature of correlations in gapped quantum many-body systems together with the finite velocity of the spread of correlations, we showed that the relevant information is already encoded in the short-time behavior of the correlation function. Furthermore, we discuss how static ground-state currents encode the same information, thereby providing a concrete experimental protocol how to extract transport coefficients from ground-state currents alone. We confirmed our general arguments by numerical simulations of the paradigmatic Chern insulator state of fermions in a completely filled Hofstadter band.

### v2

@@ Abstract
Transport properties are central to characterizing quantum matter, yet their extraction typically requires external forcing and time-resolved measurements. In this work, we propose a scheme to access transport coefficients directly from measurements of local static ground-state currents – quantities readily accessible in quantum-engineered platforms. By exploiting the exponential decay of correlations in gapped systems and the finite velocity of correlation spreading, we demonstrate that the local Hall response of correlated insulators can be reconstructed from a small set of quasi-local current observables. We derive explicit relations connecting these static observables to a practical local Chern marker, and introduce a scalable digital protocol for measuring the required generalized currents in quantum simulators. We demonstrate the applicability of our approach through numerical studies of emblematic Chern-insulator systems, both in the non-interacting and strongly-correlated (fractional) regime. Our method extends naturally to a broad class of correlated systems, even at finite temperature, offering a broadly applicable route to probing transport in engineered quantum matter.
Introduction
Figure 1: (a) Traditional methods to access transport coefficients like the Hall conductivity often use external forcing. (b) In contrast, we propose to measure ground-state currents (red arrows) within a finite radius (gray shading) around a reference site (blue dot) to extract the local Hall response (Chern marker).
@@ Introduction
Transport measurements are essential for identifying and characterizing exotic quantum matter, including superconducting states, strange metals, and topological insulators. While transport is traditionally probed by measuring currents upon connecting reservoirs to a sample, transport coefficients can also be extracted indirectly through other observables and protocols. Examples of such approaches include the density response to a magnetic perturbation following Středa’s formula [31, 30], circular dichroism via Kramers-Kronig relations [5, 34, 24] and related spectroscopic probes [33], as well as center-of-mass drifts upon acting with an external force [23, 12, 25]. These methods have proven to be powerful to probe transport coefficients in various settings, including synthetic quantum matter of cold atoms in continuous systems [17, 2] and optical lattices [1, 4, 3, 9, 20], as well as photonic systems [36, 11].
@@
Theoretically, the conductivity can be expressed as a current-current correlation function via the Kubo formula [18], suggesting that transport could be monitored by evaluating local currents in the ground state. As further discussed below, this would a priori require monitoring these currents over time, within a region located around a reference site delimited by the correlation length. While challenging in traditional solid-state platforms, this is particularly appealing to recent quantum-engineered setups, where local currents can be measured with great accuracy [15]. Local transport properties can thus be extracted, which, for topological insulating states, provides a practical scheme to access local probes of topology, e.g. local Chern markers [7, 8, 27, 6]. At this stage, however, the measurement of temporal correlations is experimentally challenging; see Ref. [16] for a proposal based on many-body Ramsey interferometry.
@@
In this work, we introduce a scheme to extract the local Hall response (Chern marker) of a gapped many-body system, which only requires local ground-state current measurements. In contrast to more conventional approaches, this scheme entirely builds on static measurements, performed in the absence of any external perturbation. Importantly, this scheme exploits the exponential decay of correlations [14] in gapped systems, together with the finite velocity of the spread of correlations [21]. Under these conditions, the relevant information can be accessed from a power series of the time-evolved current-current correlations. As a consequence, the problem reduces to measuring equal-time current correlations within a finite region surrounding the reference site. Focusing on the transverse (Hall) conductivity, we demonstrate that the associated current–current correlator admits a substantial simplification: it can be expressed in terms of a limited set of further-range current observables. Crucially, we find that static ground-state current measurements alone suffice to accurately determine the Hall conductivity. This paradigm is summarized schematically in Fig. 1. Inspired by recent experiments in cold-atom quantum simulators [15], we provide a scalable, digital protocol to access the required further-range currents. We also present a numerical validation of our scheme, considering a non-interacting Chern insulator in the Harper-Hofstadter model, as well as a strongly-correlated fractional Chern insulator of hard-core bosons.
@@
Since the correlation structures used in our protocol are common to all gapped, topological states, we expect our protocol to apply to a broad class of systems, including systems at finite temperature. Furthermore, similar protocols can be developed for other correlation functions, giving access to properties not easily accessible in traditional solid-state platforms.
Results
@@
From the Hall conductivity to local ground-state currents. Using Kubo’s formula, the local transverse conductivity, evaluated at position $\vec{r}$, reads
@@
The single-frequency ansatz. The unequal-time correlation function contained in Eq. (4) cannot be accessed in state-of-the-art quantum simulators. For gapped states, we can circumvent this limitation by approximating the correlator $\mathcal{C}(\vec{r},t)$ with an exponentially damped oscillation,
@@
This approximation is applicable under the assumption that there is a single characteristic gap $\omega_{0}$, from the ground state to the relevant low-energy excitations. The rate $\Gamma$ then characterizes the broadening of this transition due to band dispersion and finite-size effects. This single-frequency ansatz is expected to hold in the regime of relatively flat Bloch bands, as we explicitly demonstrate below and in the Methods.
@@
This expression offers an explicit recipe to obtain the local Hall response from the three parameters entering the single-frequency ansatz in Eq. (6): $\mathcal{C}(\vec{r},0)$, $\omega_{0}$ and $\Gamma$. Since the Hall response is directly related to the Chern number [32, 22], we point out that Eq. (7) provides a simple and practical expression for a local Chern marker ${\sf Ch}(\vec{r})$.
Figure 2: (a) In a Chern insulator state, the current-current correlator $\mathcal{C}(\vec{r},t)$ defined in Eq. (4), relative to a reference site $\vec{r}$, exhibits damped oscillations, which are well captured by the ansatz in Eq. (6). (b) Upon fitting the damped oscillations, one can use the expression in Eq. (7) to obtain a local Chern marker. The result is in agreement with the Chern number ${\sf Ch}=1$ of the populated band. Here, the calculations were performed by filling the lowest band of the Harper-Hofstadter model with non-interacting fermions, at flux $\alpha=\nicefrac{{1}}{{q}}$ per plaquette.
To confirm the validity of this approximation, we performed a numerical study of a paradigmatic Chern insulator state, which is obtained by filling the lowest band of the Harper-Hofstadter model with non-interacting fermions; see also the next paragraph for details. We find that the correlator $\mathcal{C}(\vec{r},t)$ that was obtained numerically is in good agreement with the ansatz in Eq. (7); see Fig. 2(a). Here, we extracted the parameters $\Gamma$ and $\omega_{0}$ by fitting the numerical data. As discussed in detail below, these parameters match the bandwidth $\Gamma\!\approx\!W$ and cyclotron gap $\omega_{0}\!\approx\!\Omega_{c}$ of the corresponding Hofstadter bands, respectively.
Furthermore, we confirm that the local Chern marker ${\sf Ch}(\vec{r})=2\pi\sigma_{\rm H}(\vec{r})$ [Eq. (7)], extracted from the fitted curve, agrees well with the expected Chern number ${\sf Ch}=1$ of the lowest Hofstadter band in the limit of sufficiently small $\alpha$; see Fig. 2(b). Deviations at large $\alpha$ are attributed to the stronger dispersion of the Hofstadter bands in this regime, which broadens the transition.
We explicitly confirm the local nature of the Chern marker ${\sf Ch}(\vec{r})$ in the Methods, where we present the results for a junction displaying two (local and distinct) Chern insulating phases.
Unequal-time correlators and static ground-state observables. As already remarked, it is challenging to measure the unequal-time correlator $\braket{\hat{j}_{x}(\vec{r},t)\hat{j}_{y}(\vec{r}^{\prime},0)}$ defined in Eq. (4), both in quantum-simulator experiments and in numerical studies of strongly-correlated systems. And while the ansatz in Eq. (6) offers a possible simplification to the problem, it would be appealing to identify a general, practical and self-contained method to systematically determine such unequal-time correlators without any a priori knowledge of the system’s characteristics (e.g. the spectral gap $\omega_{0}$). To do so, we proceed by rewriting the formal expression for this correlator in a more appealing form, using the Baker-Campbell-Hausdorff identity,
@@
Importantly, we remark that $c_{1}$ scales with the damping $\Gamma$, hence it is small whenever band-dispersion effects are limited. In fact, in our numerical investigations detailed below, we find that $c_{1}$ is indeed negligible, as well as higher odd corrections ($c_{3,5,\ldots}$). In contrast, we find that including even corrections ($c_{4,6,\ldots}$) improves the determination of the correlator $\mathcal{C}(\vec{r},t)$; see the Methods. However, including those higher-order corrections leads to a significant overhead due to the numerous currents that need to be evaluated; see below.
Furthermore, we remind that the coefficient $\omega_{0}$, which requires the measurement of $c_{2}$, corresponds to the characteristic gap in the flat-band regime. Thus, this coefficient could be evaluated independently through conventional spectroscopy. Altogether, this indicates that the measurement of the transport coefficient ${\sf Ch}(\vec{r})=2\pi\sigma_{\rm H}(\vec{r})$ can be limited to the evaluation of the single coefficient $c_{0}$ in the flat-band regime (i.e. the low-flux limit in the Harper-Hofstadter model).
Since the explicit form of the coefficients $c_{m}$ depends on the details of the Hamiltonian, we will now investigate concrete examples in the next paragraph, based on the Harper-Hofstadter model. However, we expect our approach to apply to a broad class of gapped, local Hamiltonians.
Extracting the $c_{m}$ coefficients in the Harper-Hofstadter model. We consider the Harper-Hofstadter model, defined on a square lattice, with flux $\alpha$ per plaquette. In the Landau gauge, the Hamiltonian reads
@@
where $\vec{r}=(x,y)$ denote the lattice sites, $J$ is the hopping amplitude, and $\hat{c}^{(\dagger)}_{x,y}$ are annihilation (creation) operators for fermions or bosons, depending on the context below. This well-known model exhibits topological bands [32], with Chern number ${\sf Ch}=1$ in the lowest band for $\alpha\!=\!1/q$, with $q$ an integer. The continuum limit of the model is reached upon setting $\alpha\rightarrow 0$, where the Hofstadter bands reduce to (flat) Landau levels.
@@
$\hat{j}_{\vec{\mu}}(\vec{r};\theta)=iJ\mathrm{e}^{i\theta}\hat{c}^{\dagger}_{\vec{r}+\vec{\mu}}\hat{c}^{\vphantom{\dagger}}_{\vec{r}}+\mathrm{H.c.},$ (15)
@@
We note that the number of summands in Eq. (19) is given by $8^{m}$ [Eq. (16)], with each summand contributing $4$ generalized currents [Eq. (18)]. Therefore, at $m$-th order $4\times 8^{m}$ currents need to be measured, in total resulting in $4\times(8^{0}+8^{1}+8^{2})=292$ current measurements to access the local Hall conductivity from the leading-order coefficients $c_{0,1,2}$. In Fig. 3(a-c) we show the currents that have to be evaluated for a Harper-Hofstadter model on a $13\times 13$ square lattice with the reference site chosen at the center of the system.
Figure 3: (a-c) Currents $\hat{j}_{\vec{\mu}}(\vec{R},\Theta)$ to be evaluated at (a) zeroth, (b) first, and (c) second order in $t$ for a Harper-Hofstadter model on a square lattice of $13\times 13$ sites with the reference site (blue) chosen in the center of the system. Darker arrows indicate that more than one generalized current with different phases $\Theta$ have to be measured. The necessary currents are found by first applying the iterative construction in Eq. (16), then extracting the imaginary part of the correlator using Eq. (18), and integrating over all lattice sites $\vec{R}$. (d) Local Chern marker ${\sf Ch}(\vec{r})=2\pi\sigma_{\rm H}(\vec{r})$ [Eq. (7)] as obtained from the coefficients $c_{m}$ in Eq. (10) using the experimental protocol. Here, the coefficients $c_{m}$ are connected to the parameters of Eq. (6) via Eq. (12). The values for the Chern marker obtained in this manner are in good agreement among each other as well as with the expected ${\sf Ch}=1$ for the lowest Hofstadter band. For additional details see the Methods section.
@@
We close this paragraph by reminding that the measurement of the transport coefficient ${\sf Ch}(\vec{r})=2\pi\sigma_{\rm H}(\vec{r})$ can be limited to the evaluation of the single coefficient $c_{0}$, in the flat-band regime; see above. This measurement would only require the determination of the four local diagonal currents displayed in Fig. 3(a), which is experimentally appealing.
@@
Numerical study I: Non-interacting Chern insulator. We demonstrate the applicability of our scheme through a numerical study of an emblematic Chern-insulator system, obtained by considering non-interacting fermions in the Harper-Hofstadter model. In our simulations, we consider an open square of size $L\times L=13\times 13$ with varying flux $\alpha=\nicefrac{{1}}{{q}}$, $q=3,\ldots,15$, per plaquette, so that the lowest band has Chern number ${\sf Ch}=1$. We construct a Chern insulator by occupying the lowest $N=\nicefrac{{(L_{x}\times L_{y})}}{{q}}$ states with fermions and use this state to test our experimental protocol. In particular, we implement the recursive generation of the currents visualized in Fig. 3(a-c) and evaluate their expectation values to obtain $c_{0,1,2}$. From these, we can extract the parameters of the damped oscillation [Fig. 2(a)] through Eq. (12) and insert them into the expression for the local Chern marker ${\sf Ch}(\vec{r})$ in Eq. (7).
@@
We find that the results obtained in this manner are in good agreement with the expected Chern number ${\sf Ch}=1$ of the lowest Hofstadter band, see Fig. 3(d). We emphasize again that, in this approach, the problem reduced to the evaluation of $292$ generalized ground-state currents, instead of performing explicit time-evolution. While including the higher order contribution $c_{4}$ further improves the results, we find that neglecting the odd orders ($c_{1}$ and $c_{3}$) does not affect the results; see also the Methods.
Before turning to a second application of our method, we comment on the applicability of the single-frequency ansatz in Eq. (6). We compare the characteristic oscillation frequency $\omega_{0}$, extracted from a fit of the correlator $\mathcal{C}(\vec{r},t)$ using Eq. (6), to the cyclotron gap $\Omega_{c}$. The latter quantity corresponds to the bandgap to the first excited band in the Harper-Hofstadter model, and reaches the value $\Omega_{c}^{\rm cont}=4\pi\alpha$ in the low-flux (continuum) limit. As shown in Fig. 4, the extracted oscillation frequency $\omega_{0}$ agrees well with the cyclotron frequency $\Omega_{c}$ over a broad flux window, $\alpha\lesssim 0.2$, hence validating the single-frequency ansatz in Eq. (6) in this range. We also show that the parameter $\omega_{0}$ extracted from the coefficients $c_{m}$ [see Eq. (12)] agrees with the single-frequency ansatz [Eq. (6)] throughout the entire flux range. The identification of the cyclotron bandgap as the relevant frequency indicates that the parameter $\omega_{0}$ can be determined independently through spectroscopy.
Numerical study II: Laughlin state of hard-core bosons. We complete our numerical validation by exploring the applicability of our method to a bosonic Laughlin state [19]: a strongly-correlated topological state. Specifically, we consider the Harper-Hofstadter model discussed above, and add repulsive Hubbard interactions of strength $U>0$,
For simplicity, we consider the hard-core bosonic limit ($U/J\to\infty$), where the exact expressions for the commutators in Eq. (16) remain unchanged. At magnetic filling factor $\nu=\nicefrac{{1}}{{2}}$, it is well-established that this model hosts a lattice analog of the corresponding Laughlin state [28].
For large systems, while ground-state searches and evaluations of (quasi-)local currents are numerically feasible, long-time evolution of the strongly-correlated many-body system becomes increasingly challenging. Therefore, we apply the protocol outlined above to extract the local Chern marker from static ground-state currents. To this end, we employ matrix product states (MPS) and the density-matrix renormalization group (DMRG) to variationally find the ground state [37, 26], using the MPSKit toolkit [13].
For concreteness, we fix the flux $\alpha=\nicefrac{{1}}{{6}}$ per plaquette and vary the number of particles $N$ and the system size $L$ to realize a $\nicefrac{{1}}{{2}}$-Laughlin state. We measure the currents indicated in Fig. 3(a-c) and use them to obtain $c_{0,1,2}$ and – through Eq. (12) – the local many-body Chern marker $\mathsf{Ch}(\vec{r})$ for a reference site in the center of the system. We find that for sufficiently large systems ($N\geq 6$, $L\geq 10$) the extracted results converge and are in good agreement with the expected $\mathsf{Ch}\!=\!\nicefrac{{1}}{{2}}$ of the Laughlin state; see Fig. 5. We attribute the remaining deviations from the expected value even for large systems to band-dispersion effects due to the relatively large flux $\alpha=\nicefrac{{1}}{{6}}$; see also Fig. 3(d). We remark that the characteristic frequency $\omega_{0}$ matches well the cyclotron gap of the system; see the inset of Fig. 5. We attribute this agreement to residual Landau-level selection rules, which are expected to become exact in the low-flux limit of the Harper-Hofstadter model; see also Ref. [35].
Our findings indicate that, despite the simplifying assumptions and strong correlations of the Laughlin state, our approach is capable of efficiently extracting a local topological marker with a satisfactory level of accuracy. This is due to the finite correlation length present in the Laughlin state, originating from the many-body bulk gap. This key property enables the establishment of a local Chern marker in strongly-correlated topological states, as similarly demonstrated by the local Středa response [25].
Figure 5: Local many-body Chern marker $\mathsf{Ch}(\vec{r})$ [Eq. (7)] for a reference site in the center of an $L\times L$-system of $N$ hard-core bosons at $\alpha=\nicefrac{{1}}{{6}}$ in the Laughlin regime. The marker is obtained via Eq. (12) by evaluating the coefficients $c_{0},c_{1},$ and $c_{2}$ in Eq. (10) from the static ground-state. We find that for sufficiently large $N\geq 6$, the extracted Chern marker is in good agreement with the expected $\sf{Ch}=\nicefrac{{1}}{{2}}$ of the Laughlin state (dashed line). The inset shows the extracted frequency $\omega_{0}$, which is in good agreement with the cyclotron frequency of the non-interacting model for large NN (dashed line).
@@
Protocol for measuring further-range currents. The experimental protocol presented above requires the measurement of generalized, further-range currents connecting lattice sites $\vec{R}$ and $\vec{R}+\vec{\mu}$. Here, we present a scalable, digital protocol to achieve such measurements using only nearest-neighbor tunnelings $\hat{T}_{ij}$ and local energy offsets $\hat{M}_{i}$ on the lattice,
@@
as well as local density measurements on the sites involved.
@@
  • $\ell=3$: Here, we need a pulse of duration $t_{2}=\frac{\pi}{2\Delta}$ to obtain the required phase (up to a global sign). The initial ground-state densities contribute with $\gamma_{1}\!=\!\gamma_{3}\!=\!\frac{1}{2}$ and $\gamma_{2}\!=\!-1$ along the path; see Fig. 6.
@@
Figure 6: Upper panel: Sketch of the pulse sequence needed to evaluate further-range currents. Lower left: Ground state density (dots) and $m=0$ currents (red arrows) of the Chern insulator in the lowest band of a Harper-Hofstadter model at $\alpha=\nicefrac{{1}}{{5}}$. The blue arrow indicates a possible path to measure a current between next-nearest neighbors. Lower right: Density difference after the pulse sequence and coefficients $\gamma_{i}$ with which the respective points along the path are weighted.
Discussion
In this work, we showed that measurements of ground-state currents of gapped many-body systems are sufficient to extract the local Hall marker. In particular, using the quasi-local nature of correlations in gapped quantum many-body systems together with the finite velocity of the spread of correlations, we showed that the relevant information is already encoded in the short-time behavior of the correlation function. Furthermore, we discuss how static ground-state currents encode the same information, thereby providing a concrete experimental protocol to extract transport coefficients from ground-state currents alone. We confirmed our general arguments by numerical simulations of the paradigmatic Chern insulator state of non-interacting fermions in a completely filled Hofstadter band, as well as the strongly-correlated bosonic Laughlin state in a half-filled Hofstadter band.